OpenStudy (anonymous):

sin Ɵ / 1 + sin Ɵ - sin Ɵ / 1 - sin Ɵ Complete the identity ..

OpenStudy (luigi0210):

\[\frac{ \sin \theta }{ 1+\sin \theta }-\frac{ \sin \theta }{ 1-\sin \theta }\]

OpenStudy (luigi0210):

Is this your problem?

OpenStudy (anonymous):

Yeah !

OpenStudy (luigi0210):

Okay, first what we do is get common denominators. Can you do that?

OpenStudy (anonymous):

\[\cos^2 \theta \] ?

OpenStudy (luigi0210):

Close enough: 1-sin^2x

OpenStudy (anonymous):

wait how did you get that?

OpenStudy (shamim):

\[\frac{ \sin \theta (1- \sin \theta)-\sin \theta(1+\sin \theta) }{ (1-\sin \theta)(1+\sin \theta) }\]

OpenStudy (luigi0210):

@shamim illustrates it, thank you for that!

OpenStudy (shamim):

\[\frac{ \sin \theta (1-\sin \theta-1-\sin \theta) }{ 1-\sin ^{2}\theta }\]

OpenStudy (anonymous):

I'm so confused....

OpenStudy (luigi0210):

@shamim can take care of you from here. Call if you need anything!

OpenStudy (shamim):

where u confused

OpenStudy (anonymous):

What is the first step ?

OpenStudy (shamim):

\[\frac{ -2 \sin ^{2}\theta }{ \cos ^{2}\theta }\]

OpenStudy (anonymous):

ohh ok so you get -2tan^2 theta but i dont know how to get to there

OpenStudy (shamim):

ok

OpenStudy (shamim):

just follw my every step

OpenStudy (shamim):

go back

OpenStudy (anonymous):

how did it turn in multiplication?

OpenStudy (shamim):

u know\[(a+b)(a-b)=a ^{2}-b ^{2}\]

OpenStudy (shamim):

do u know this formula?

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

OHHHHHHHHHHHHHHH I SSEE NOW

OpenStudy (shamim):

then\[(1+ \sin \theta)(1- \sin \theta)=1^{2}-\sin ^{2}\theta\]

OpenStudy (anonymous):

\[\sin \theta(1-\sin \theta-1-\sin \theta) / 1-\sin^2 \theta \] im stuck here

OpenStudy (anonymous):

):

OpenStudy (anonymous):

where did the -2sin^2 theta come from ?

OpenStudy (luigi0210):

Did he leave?

OpenStudy (anonymous):

I think so... got tired of my dumbness :C

OpenStudy (luigi0210):

No, you're not dumb at all.

OpenStudy (luigi0210):

But I can try to finish it if you want?

OpenStudy (anonymous):

Ok please thank you

OpenStudy (luigi0210):

Where did you guys leave off at?

OpenStudy (anonymous):

I dont know where to go after sinƟ(1-sinƟ-1-sinƟ) / cos^2Ɵ

OpenStudy (anonymous):

how did the top part turn in -2sin^2Ɵ ?

OpenStudy (luigi0210):

Well using HIS method: